OpenStudy (anonymous):

Which of the following is the simplified form of (4x^3)/(9y^4)÷(8x^5)/(3y^2)

OpenStudy (kropot72):

\[\frac{4x ^{3}}{9y ^{4}}\times \frac{3y ^{2}}{8x ^{5}}\] The first part of simplification is to divide numerator and denominator by 4x^3. Can you do that?

jhonyy9 (jhonyy9):

do you know how divide two fractions ?

jhonyy9 (jhonyy9):

you need cooperat anything ok ?

jhonyy9 (jhonyy9):

come on please i like to help you understanding

jhonyy9 (jhonyy9):

ok so when you need divide two fractions you need multiplie first by second inversed ok ? how you have wrote there little up kropot72 ok ?

jhonyy9 (jhonyy9):

@kropot72 please end your work

OpenStudy (kropot72):

When you divide numerator and denominator by 4x^3 you get the following: \[\frac{3y ^{2}}{9y ^{4}\times 2x ^{2}}\] Next divide numerator and denominator by3y^2 and get the following: \[\frac{1}{3y ^{2}\times 2x ^{2}}=\frac{1}{6x ^{2}y ^{2}}\]